Two flat parallel metal plates separated by d 20.0 cm have been given equal and opposite charges, which results in a more or less uniform electric field between them. A proton (mass m 1.67 x 10-27 kg) is released from rest at the positively charged plate A, and then accelerates uniformly until it arrives at the negatively charged plate B with a speed of 4.00 x 10 m/s.B Ignore gravity. a) Find the acceleration of the proton, and from that determine the magni plates. Be sure to include correct units itude of the electric field E between the b) Using the concept of conservation of energy, determine the change in electric potential energy APE of the proton (in joules) as it moved from A to B c) Use APE to find the difference in electric potential between the plates, VAB VA-VB, in volts. Be careful with +/- signs. Which plate is at the higher electrie potential (or voltage)? d) If instead a sodium ion Na+ (mas 3.84 x 10-26 kg) is released from rest at plate A and accelerates to plate B. what dange in electric potential energy ΔΡΕ will occur?
Answer #1

a)We know that

v^2 = u^2 + 2 a s

a = (v^2 - u^2)/2 s

a = [(4 x 10^5)^2 - 0^2]/(2 x 0.2) = 4 x 10^11 m/s^2

Hence, a = 4 x 10^11 m/s^2

We know that,

F = ma = qE

E = ma/q

E = 1.67 x 10^-27 x 4 x 10^11/(1.6 x 10^-19) = 4.18 x 10^3 N/C

Hence, E = 4.18 x 10^3 N/C

b)Change in elecrtic PE

Del-PE = 1/2 m (v^2 - u^2)

Del-PE = 0.5 x 1.67 x 10^-27 [(4 x 10^5)^2 - 0^2] = 1.34 x 10^-16 J

Hence, Del-PE = 1.34 x 10^-16 J

c)del-PE = q del-V

del-V = del-PE/q

del-V = Va - Vb = 1.34 x 10^-16/(1.6 x 10^-19) = 837.5 V

Hence, Va - Vb = 837.5 V

d)In this case also

a = 4 x 10^11 m/s^2

Del-PE = 1/2 m (v^2 - u^2)

Del-PE = 0.5 x 3.84 x 10^-26 [(4 x 10^5)^2 - 0^2] = 3.072 x 10^-15 J

Hence, del-PE = 3.072 x 10^-15 J